Grade 11Physical Chemistry

500ml of 0.1N solution of AgNO3 is added to 500ml of 0.1N Kcl solution. Calculate the concentration of nitrate in the resulting mixture.

Adhiraj Kapoor
9 Years agoGrade 11

1 Answer

Himanshu
9 Years ago
Gram equivalents of NO3 in AgNO3 = 0.500 × 0.1 = 0.05 gmeq
There is no gmeqs of NO3 in KCl.
Total volume = 500ml + 500ml = 1L
Hence, concentration of NO3 = 0.05 / 1 = 0.05N